# SUBA-Jet: a new Model for Jets in Heavy Ion Collisions

### v1

@@ Appendix D Calibration
The calibration to the IOE spectrum (see Fig. 43) reveals that the optimal choice for the gluon phase accumulation criteria is $\varphi_{c}\approx 6.0$. For $T=400$ MeV and $\alpha_{s}=0.3$, the optimal prescription for $m_{g}^{\text{reg}}$ is found to be $3.2$ MeV. Following eq. (140), one thus takes
$m_{g}^{\rm reg}(k^{+}\gg T)\approx 0.008\,\sqrt{\frac{\alpha_{s}}{0.3}}\times T\,.$ (144)
@@
Fig. 43 shows the comparison between our MC and the IOE spectrum for our specified parameters and with the optimal prescription for $\varphi_{c}$ and $m_{g}^{\text{reg}}$. The energy-weighted spectrum $\omega\,\mathrm{d}N/\mathrm{d}\omega$ is shown, to better illustrate the transition from the LPM regime to the fully coherent regime, happening around $\omega\sim\omega_{c}\equiv\hat{q}_{0}L^{2}/2\approx 48$ GeV, where $\hat{q}_{0}$ is the transport coefficient stripped of its Coulomb logarithm. Both our MC and the IOE reproduce the BDMPS-Z expectation in the LPM regime and the GLV result in the fully coherent limit. The agreement between our MC and the IOE is fulfilled nicely in the regions where the IOE is expected to be valid. For both very low (Bethe-Heitler) and very large ($z\to 1$) energies, the IOE is not expected to hold. Around the transition region, $\omega\sim\omega_{c}$, the IOE spectrum exhibits a sharp transition, which seems unphysical. On the other hand, our MC shows a smoother transition.

### v2

@@ Appendix C Calibration
The calibration to the IOE spectrum (see Fig. 50) reveals that the optimal choice for the gluon phase accumulation criteria is $\varphi_{c}\approx 6.0$. For $T=400$ MeV and $\alpha_{s}=0.3$, the optimal prescription for $m_{g}^{\text{reg}}$ is found to be $56$ MeV. Following eq. (148), one thus takes
$m_{g}^{\rm reg}(k^{+}\gg T)\approx 0.14\,\sqrt{\frac{\alpha_{s}}{0.3}}\times T\,.$ (152)
@@
Fig. 50 shows the comparison between our MC and the IOE spectrum for our specified parameters and with the optimal prescription for $\varphi_{c}$ and $m_{g}^{\text{reg}}$. The energy-weighted spectrum $\omega\,\mathrm{d}N/\mathrm{d}\omega$ is shown, to better illustrate the transition from the LPM regime to the fully coherent regime, happening around $\omega\sim\omega_{c}\equiv\hat{q}_{0}L^{2}/2\approx 196$ GeV, where $\hat{q}_{0}$ is the transport coefficient stripped of its Coulomb logarithm. Both our MC and the IOE reproduce the BDMPS-Z expectation in the LPM regime and the GLV result in the fully coherent limit. The agreement between our MC and the IOE is fulfilled nicely in the regions where the IOE is expected to be valid. For both very low (Bethe-Heitler) and very large ($z\to 1$) energies, the IOE is not expected to hold. Around the transition region, $\omega\sim\omega_{c}$, the IOE spectrum exhibits a sharp transition, which seems unphysical. On the other hand, our MC shows a smoother transition.